Can someone please help me understand how we calcu Can someone please help me understand how we calculate and answer #5. I already got number #4 and that information is used to answer #5.

Thank you so much!
Answer #1

frequencyf = 1/2pisqrt(k/m)

f = 200 N

f = mg

m = 200/10 = 20 kg

f = 1/2*3.14*sqrt(2000/20)

f = 1.59 Hz